### v1

@@ Abstract
keywords:
Large Language Models, LLM Agents, Financial Analysis, Stock Selection, SEC filings, Retrieval Augmented Generation, MarketSenseAI
pacs:
pacs:
JEL Classification]C45, C61, G11, G15
MSC Classification]68T07, 68T50, 91G10, 91G15
1 Introduction
@@
    Comprehensive Real-World Evaluation: Experiments using S&P 100 stocks for a two-year period (2023–2024) and S&P 500 stocks for 2024 illustrate the robustness of our proposed system, revealing a notable improvement in fundamental analysis accuracy and consistent excess returns of 8.0–18.9% with comparable risk over benchmark indices.
@@
The remainder of this paper is structured as follows: Section 2 provides a literature review examining current research in LLM-based systems for financial analysis. Section 3 details updates to the MarketSenseAI architecture, including agent responsibilities and data flow. Section 4 presents our experimental design, covering datasets, evaluation metrics, and baseline comparisons. Section 5 discusses empirical findings from S&P 100 and S&P 500 stocks, including performance metrics, risk-adjusted returns, and a factor analysis. Finally, Section 6 concludes with key insights and outlines future developments for MarketSenseAI.
2 Background and Related Work
@@
2.1 LLM-Based Fundamental Analysis
@@
2.2 Advanced Methods in LLM-Driven Investment Analysis
@@
2.3 Retrieval-Augmented Techniques
@@
2.4 Importance of Filings and Earnings Calls in Fundamental Research
@@
2.5 Macroeconomic environment impact in stock analysis
@@
3 Methods
3.1 Overview of MarketSenseAI components
@@
Figure 1: Conceptual Architecture of MarketSenseAI, highlighting the core components/agents, data flow, and outcomes of each agent for a selected stock (i.e., Nvidia on Jan. 3, 2025). The agents’ outputs have been condensed for illustration purposes.
@@
    Fundamentals Agent: Focuses on analyzing each company’s financial statements (e.g., balance sheets, income statements, and cash flow reports). To handle large and often complex numerical data, these statements are preprocessed and reduced into abbreviated formats (e.g., grouping figures in “million” or “billion”) before the LLM extracts key insights. The system compares recent quarters to highlight shifts in profitability, revenue, or leverage ratios, laying the groundwork for fundamental analysis. Besides the numerical figures, the updated
@@
    Signal Agent: The final component integrates the textual outputs from the previous four modules—news, fundamentals, price dynamics, and macroeconomic outlook—into a single decision-making process. Implemented via a Chain-of-Thought (CoT) prompting strategy, the LLM reviews each aggregated summary to produce an investment signal (buy, hold, or sell). It also provides a written explanation that traces the reasoning behind each recommendation, thereby enhancing transparency and interpretability.
@@
3.2 Enhanced fundamentals analysis
@@
3.2.1 A Three-Layer Approach to Integrating Qualitative and Quantitative Data
@@
3.2.2 Evaluating the Impact of SEC Filings and Earnings Calls
To assess how SEC filings and earnings call data affect the Fundamentals Agent’s outputs, we conducted a sentiment analysis on 1,500 generated summaries covering S&P500 stocks at three different points in time. The FinBERT model was utilized to obtain the sentiment of each generated summary [46]. Table1 and Fig. 3 reveal distinct patterns between outputs with and without this additional text-based information. When incorporating filings and calls data, the analysis showed a slightly less positive average sentiment (Mean = 0.31) with more moderate variance (Std Dev = 0.28). In contrast, analyses based on numerical data alone exhibited more positively skewed results (Mean = 0.36) with a wider spread of sentiment values (Std Dev = 0.40). This moderation in sentiment when including filings data is particularly noteworthy as SEC filings require companies to disclosure risks and uncertainties in dedicated sections, even when their financial metrics appear strong, thus providing a more balanced perspective of the company’s outlook.
@@ 3.2.2 Evaluating the Impact of SEC Filings and Earnings Calls
Table 1: Statistics of sentiment analysis of Fundamentals Agent’s output across stocks and dates.
Statistic Sentiment (Full) [^1] Full includes SEC Filings and Earnings Call transcripts Sentiment (Basic) [^2] Basic includes only numerical data from quarterly statements Difference
@@
[…]
@@
3.3 Macroeconomic Analysis Improvements
@@
3.3.1 Data Injection
The data injection stage (Fig. 5) is designed to efficiently collect, process, and store macroeconomic reports from multiple sources, including central banks (e.g., Federal Reserve, European Central Bank), statistical bureaus, the International Monetary Fund (IMF), the Bank for International Settlements (BIS), and opinionated expert briefings from global investment banks such as JPMorgan and BlackRock. We have implemented institution-specific parsing scripts that handle the unique formatting and structure of reports from each source, ensuring consistent and accurate data extraction across different providers.
@@ 3.3.1 Data Injection
Metadata Extraction and Filtering
Once a document is parsed, we identify key attributes like publication date, publisher, and URL. These metadata not only ensure document provenance but also enables the system to sequence reports chronologically. Next, an LLM-powered classifier determines whether the text is relevant to macroeconomic analysis. Irrelevant documents (e.g., marketing brochures) are discarded at this step.
Content Cleaning and Summarization
For relevant documents, another LLM process removes extraneous text (e.g., disclaimers, duplicate headers) and produces a summary capturing the document’s core insights. Large files (over 30 pages) are broken into smaller chunks; each chunk is cleaned, summarized, and then consolidated into a single refined representation of the entire document. This approach preserves vital macroeconomic details without overwhelming LLM context limits.
Storage and Indexing
The cleaned content, along with metadata, is stored. Parallelly, a look up table is updated with relevant metadata to maintain an organized inventory of all processed documents. Afterward, we conduct semantic chunking of new reports [47]; each chunk is embedded and stored in a Vector Datastore for fast, similarity-based retrieval. By chunking on natural boundaries (e.g., the end of a section or a shift in economic theme), the system ensures granular and semantically coherent indexing of macroeconomic information [48].
3.3.2 Macroeconomic Data Generation
@@
  • MarketSenseAI (Predefined Queries & HyDE): For single-stock analysis, we employ a HyDE approach on a fixed set of queries (e.g., “U.S. macro outlook,” “investment opportunities and risks”). This yields brief, rounded macroeconomic insights without overburdening the final signal-generation stage. An example of this output is given at Table 2.
@@
Table 2: Macroeconomic Agent’s Output (January 3, 2025) used in MarketSenseAI
@@
3.3.3 Retrieval Performance Evaluation
To assess the retrieval pipeline’s ability to handle macroeconomic queries of varying complexity, we tested three methods (Simple, Optimized, and HyDE) across different chunk sizes, evaluating context recall, context precision, answer relevancy, and faithfulness [49]. Each approach shapes how queries are transformed before performing semantic similarity searches in the vector database, thereby influencing which top-$n$ chunks are retrieved. The results in Table 3 demonstrate the effectiveness of different retrieval methods across varying chunk sizes for complex macroeconomic queries. These findings offer several key insights:
  • Context Precision remains high ($\geq 0.98$) in all configurations, indicating that even when queries span multiple reports, irrelevant chunks are not in the top-$n$ results. This supports the validity of the data injection presented in Section 3.3.1.
@@ 3.3.3 Retrieval Performance Evaluation
Table 3: Performance Comparison of Retrieval Methods by Chunk Size
@@ 4.1 Data
4.2 Technology Stack
The GPT-4o model serves as the primary LLM for all processes requiring model inference [50], while the system maintains an LLM-agnostic architecture that allows seamless integration of alternative models via API. For portfolio analysis and strategy validation, we utilized VectorBTPro [^2] https://vectorbt.pro/, which provided robust tools for backtesting financial strategies while accounting from transaction costs. To assess the RAG methods outlined in Section 3.3, we employed the Ragas framework [51], leveraging GPT-4o-mini for cost efficiency. While this choice may have impacted evaluation results compared to the full-scale GPT-4o model, this did not affect the relative comparison of the methods under evaluation.
The vector datastore is based on Pinecone [^3] https://www.pinecone.io/ and the agents within the system are built on OpenAI’s client. The RAG processes leverage the LlamaIndex [^4] https://www.llamaindex.ai/ framework, enabling effective retrieval and augmented generation workflows.
For data collection, macroeconomic reports are scraped using tools like Selenium and BeautifulSoup, combined with custom scripts tailored to specific data sources. SEC filings are sourced directly from the SEC’s EDGAR API, while earnings call transcripts are obtained via RapidAPI [^5] https://rapidapi.com/, which aggregates data from platforms such as SeekingAlpha and MarketBeat.
4.3 Evaluation Approach
In addition to the agent-specific evaluation presented in Section 3, we evaluate the quality of MarketSenseAI’s signals by constructing portfolios and comparing their performance against relevant benchmarks. Specifically, we focus on long-only portfolios based on MarketSenseAI’s buy signals, implemented in two forms: equally weighted and market capitalization weighted. These portfolios are compared against their corresponding equally or market cap weighted benchmark (S&P 100 or S&P 500) to assess the system’s effectiveness in generating actionable investment signals. The evaluated signals/strategies and their relevant benchmarks are presented in Table 4. Typical performance and risk metrics were used for assessing both the MarketSenseAI-based and the benchmark portfolios that are described at Table 5.
Table 4: Investment Strategies and Benchmark Portfolios
@@ 4.3 Evaluation Approach
[…]
Table 5: Portfolio Evaluation Metrics
Metric Description
Total Return The portfolio cumulative returns (%) over a specific period
Sharpe Ratio A measure of risk-adjusted return; calculated as the average return earned in excess of the risk-free rate per unit of volatility
Sortino Ratio Similar to the Sharpe Ratio, but measures returns relative to downside risk, focusing on negative asset volatility
Volatility (Vol) A statistical measure of the dispersion of returns for a given security or market index, measured using standard deviation
Win Rate (Win%) The percentage of trades that are profitable out of the total number executed
Maximum Drawdown (MDD) The maximum observed percentage loss from a peak to a trough of a portfolio, before a new peak is attained
[…]
5 Evaluation Results
@@
5.1 Overall Performance Overview
MarketSenseAI’s ability to identify outperforming equities is evident across multiple dimensions. In the S&P 100 universe, the system’s selected stocks achieved a 125.9% cumulative return under market cap-weighting (MS-Cap), significantly surpassing the S&P 100 index return of 73.5% (Table 6). This outperformance persisted in equal-weighted portfolios (MS-Eq), where selected equities returned 55.7% versus 42.3% for the equal-weighted S&P 100. Critically, these gains were not achieved through excessive risk-taking: the MS-Cap portfolio exhibited a 16% higher Sortino ratio (4.43 vs. 3.82) compared to the cap-weighted benchmark, despite experiencing higher volatility.
Table 6: Performance Metrics (2023-2024)
Portfolio Return [^1] Values in parentheses represent the total returns (%) after transaction costs (10bps/trade). Sharpe Sortino Vol MDD MDDd [^2] The duration of Maximum Drawdown (MDD) in days.
@@ 5.1 Overall Performance Overview
[…]
The system’s selection capability scaled effectively with market breadth. When applied to the S&P 500 universe during 2024, MarketSenseAI’s selected equities delivered 25.8% returns in equal-weighted portfolios compared to 12.8% for the S&P 500 Equal Weight benchmark, representing a 102% relative outperformance. This expansion to a broader universe also improved risk-adjusted performance, with the Sortino ratio increasing from 3.25 (S&P 100 MS-Eq) to 3.68 (S&P 500 MS-Eq). Alpha generation improved correspondingly, rising from 8.0% in the S&P 100 MS-Eq to 18.9% in the S&P 500 MS-Eq (Table 7), confirming the system’s enhanced ability to identify opportunities in larger universes.
Table 7: Performance Attribution Analysis
Portfolio Beta Alpha (%) Total Trades Win Rate (%) Buy Signals [^1] Values in the parentheses represent the standard deviation of the average number of buy signals per month.

### v2

@@ Abstract
I Introduction
@@
    Detailed Real-World Evaluation: Experiments using S&P 100 stocks for a two-year period (2023–2024) and S&P 500 stocks for 2024 illustrate the robustness of our proposed system, revealing a notable improvement in fundamental analysis accuracy and consistent excess returns of 8.0–18.9% with comparable risk over benchmark indices.
@@
The remainder of this paper is structured as follows: Section II provides a literature review examining current research in LLM-based systems for financial analysis. Section III details updates to the MarketSenseAI architecture, including agent responsibilities and data flow. Section IV presents our experimental design, covering datasets, evaluation metrics, and baseline comparisons. Section V discusses empirical findings from S&P 100 and S&P 500 stocks, including performance metrics, risk-adjusted returns, and a factor analysis. Finally, Section VI concludes with key insights and outlines future developments for MarketSenseAI.
II Background and Related Work
@@
II-A LLM-Based Fundamental Analysis
@@
II-B Advanced Methods in LLM-Driven Investment Analysis
@@
II-C Retrieval-Augmented Techniques
@@
II-D Importance of Filings and Earnings Calls in Fundamental Research
@@
II-E Macroeconomic environment impact in stock analysis
@@
III Methods
III-A Overview of MarketSenseAI components
Figure 1: Conceptual Architecture of MarketSenseAI, highlighting for a selected stock (i.e., Nvidia on Jan. 3, 2025). The agents’ outputs have been condensed for illustration purposes.
@@ III-A Overview of MarketSenseAI components
    Fundamentals Agent: Focuses on analyzing each company’s financial statements (e.g., balance sheets, income statements, and cash flow reports). To handle large and often complex numerical data, these statements are preprocessed and reduced into abbreviated formats (e.g., grouping figures in “million” or “billion”) before the LLM extracts key insights. The system compares recent quarters to highlight shifts in profitability, revenue, or leverage ratios, laying the groundwork for fundamental analysis. Besides the numerical figures, the updated agent analyzes SEC filings and earnings call transcripts as analyzed at Section III-B.
@@
    Signal Agent: The final component integrates the textual outputs from the previous four modules—news, fundamentals, price dynamics, and macroeconomic outlook—into a single decision-making process. Implemented via a CoT prompting strategy, the LLM reviews each aggregated summary to produce an investment signal (buy, hold, or sell). It also provides a written explanation that traces the reasoning behind each recommendation, thereby enhancing transparency and interpretability.
@@
III-B Enhanced fundamentals analysis
@@
III-B1 A Three-Layer Approach to Integrating Qualitative and Quantitative Data
@@
III-B2 Evaluating the Impact of SEC Filings and Earnings Calls
To assess how SEC filings and earnings call data affect the Fundamentals Agent’s outputs, we conducted a sentiment analysis on 1,500 generated summaries covering S&P500 stocks at three different points in time. The FinBERT model was utilized to obtain the sentiment of each generated summary [46]. TableI and Fig. 3 reveal distinct patterns between outputs with and without this additional text-based information. When incorporating filings and calls data, the analysis showed a slightly less positive average sentiment (Mean = 0.31) with more moderate variance (Std Dev = 0.28). In contrast, analyses based on numerical data alone exhibited more positively skewed results (Mean = 0.36) with a wider spread of sentiment values (Std Dev = 0.40). This moderation in sentiment when including filings data is particularly noteworthy as SEC filings require companies to disclosure risks and uncertainties in dedicated sections, even when their financial metrics appear strong, thus providing a more balanced perspective of the company’s outlook.
@@ III-B2 Evaluating the Impact of SEC Filings and Earnings Calls
TABLE I: Statistics of sentiment analysis of Fundamentals Agent’s output across stocks and dates.
Statistic Sentiment (Full)a Sentiment (Basic)b Difference
@@
aFull includes SEC Filings and Earnings Call transcripts.
bBasic includes only numerical data from quarterly statements.
@@
III-C Macroeconomic Analysis Improvements
@@
III-C1 Data Injection
The data injection stage (Fig. 5) is designed to efficiently collect, process, and store macroeconomic reports from multiple sources, including central banks (e.g., FED, ECB), statistical bureaus, the International Monetary Fund, the Bank for International Settlements, and sell-side reports from global investment banks such as JPMorgan and BlackRock. We have implemented institution-specific parsing scripts that handle the unique formatting and structure of reports from each source, ensuring consistent and accurate data extraction across different providers.
@@ III-C1 Data Injection
Metadata Extraction and Filtering: Once a document is parsed, we identify key attributes like publication date, publisher, and URL. These metadata not only ensure document provenance but also enables the system to sequence reports chronologically. Next, an LLM-powered classifier determines whether the text is relevant to macroeconomic analysis. Irrelevant documents (e.g., marketing brochures) are discarded at this step.
Content Cleaning and Summarization: For relevant documents, another LLM process removes extraneous text (e.g., disclaimers, duplicate headers) and produces a summary capturing the document’s core insights. Large files (over 30 pages) are broken into smaller chunks; each chunk is cleaned, summarized, and then consolidated into a single refined representation of the entire document. This approach preserves vital macroeconomic details without overwhelming LLM context limits.
Storage and Indexing: The cleaned content, along with metadata, is stored. Parallelly, a look up table is updated with relevant metadata to maintain an organized inventory of all processed documents. Afterward, we conduct semantic chunking of new reports [47]; each chunk is embedded and stored in a Vector Datastore for fast, similarity-based retrieval. By chunking on natural boundaries (e.g., the end of a section or a shift in economic theme), the system ensures granular and semantically coherent indexing of macroeconomic information [48].
III-C2 Macroeconomic Data Generation
@@
  • MarketSenseAI (Predefined Queries & HyDE): For single-stock analysis, we employ a HyDE approach on a fixed set of queries (e.g., “U.S. macro outlook,” “investment opportunities and risks”). This yields brief, rounded macroeconomic insights without overburdening the final signal-generation stage. An example of this output is given at Table II.
@@
TABLE II: Macroeconomic Agent’s Output (January 3, 2025) used in MarketSenseAI
@@
III-C3 Retrieval Performance Evaluation
To assess the retrieval pipeline’s ability to handle macroeconomic queries of varying complexity, we tested three methods (Simple, Optimized, and HyDE) across different chunk sizes, evaluating context recall, context precision, answer relevancy, and faithfulness [49]. Each approach shapes how queries are transformed before performing semantic similarity searches in the vector database, thereby influencing which top-$n$ chunks are retrieved. The results in Table III demonstrate the effectiveness of different retrieval methods across varying chunk sizes for complex macroeconomic queries. These findings offer several key insights:
  • Context Precision remains high ($\geq 0.98$) in all configurations, indicating that even when queries span multiple reports, irrelevant chunks are not in the top-$n$ results. This supports the validity of the data injection presented in Section III-C1.
@@ III-C3 Retrieval Performance Evaluation
TABLE III: Performance Comparison of Retrieval Methods by Chunk Size
@@ IV-A Data
IV-B Technology Stack
The GPT-4o model serves as the primary LLM for all processes requiring model inference [50], while the system maintains an LLM-agnostic architecture that allows seamless integration of alternative models via API. For portfolio analysis and strategy validation, we utilized VectorBTPro, which provided robust tools for backtesting financial strategies while accounting from transaction costs. To assess the RAG methods outlined in Section III-C, we employed the Ragas framework [51], leveraging GPT-4o-mini for cost efficiency. While this choice may have impacted evaluation results compared to the full-scale GPT-4o model, this did not affect the relative comparison of the methods under evaluation.
The vector datastore is based on Pinecone and the agents within the system are built on OpenAI’s client. The RAG processes leverage the LlamaIndex framework.
For data collection, macroeconomic reports are scraped using tools like Playwrite, combined with custom scripts tailored to specific data sources. SEC filings are sourced directly from the SEC’s EDGAR API, while earnings call transcripts are obtained via RapidAPI, which aggregates data from platforms such as SeekingAlpha and MarketBeat.
IV-C Evaluation Approach
In addition to the agent-specific evaluation presented in Section III, we evaluate the quality of MarketSenseAI’s signals by constructing portfolios and comparing their performance against relevant benchmarks. Specifically, we focus on long-only portfolios based on MarketSenseAI’s buy signals, implemented in two forms: equally weighted and market capitalization weighted. These portfolios are compared against their corresponding equally or market cap weighted benchmark (S&P 100 or S&P 500) to assess the system’s effectiveness in generating actionable investment signals. The evaluated signals/strategies and their relevant benchmarks are presented in Table IV. Typical performance and risk metrics were used for assessing both the MarketSenseAI-based and the benchmark portfolios, including: total return (cumulative portfolio returns over the period), Sharpe ratio (risk-adjusted return relative to volatility), Sortino ratio (return relative to downside risk), volatility (standard deviation of returns), win rate (percentage of profitable trades), and maximum drawdown (MDD, peak-to-trough portfolio loss).
TABLE IV: Investment Strategies and Benchmark Portfolios
@@ IV-C Evaluation Approach
V Results
@@
V-A Overall Performance Overview
MarketSenseAI’s ability to identify outperforming equities is evident across multiple dimensions. In the S&P 100 universe, the system’s selected stocks achieved a 125.9% cumulative return under market cap-weighting (MS-Cap), significantly surpassing the S&P 100 index return of 73.5% (Table V). This outperformance persisted in equal-weighted portfolios (MS-Eq), where selected equities returned 55.7% versus 42.3% for the equal-weighted S&P 100. Critically, these gains were not achieved through excessive risk-taking: the MS-Cap portfolio exhibited a 16% higher Sortino ratio (4.43 vs. 3.82) compared to the cap-weighted benchmark, despite experiencing higher volatility.
TABLE V: Performance Metrics (2023-2024)
Portfolio Returna Sharpe Sortino Vol MDD MDDdb
@@ V-A Overall Performance Overview
aValues in parentheses represent the total returns (%) after transaction costs (10bps/trade).
bThe duration of Maximum Drawdown (MDD) in days.
The system’s selection capability scaled effectively with market breadth. When applied to the S&P 500 universe during 2024, MarketSenseAI’s selected equities delivered 25.8% returns in equal-weighted portfolios compared to 12.8% for the S&P 500 Equal Weight benchmark, representing a 102% relative outperformance. This expansion to a broader universe also improved risk-adjusted performance, with the Sortino ratio increasing from 3.25 (S&P 100 MS-Eq) to 3.68 (S&P 500 MS-Eq). Alpha generation improved correspondingly, rising from 8.0% in the S&P 100 MS-Eq to 18.9% in the S&P 500 MS-Eq (Table VI), confirming the system’s enhanced ability to identify opportunities in larger universes.
TABLE VI: Performance Attribution Analysis
Portfolio Beta Alpha (%) Total Trades Win Rate (%) Buy Signalsa
@@
aValues in the parentheses represent the standard deviation of the average number of buy signals per month.
Figure 7: Cumulative returns of MarketSenseAI buy monthly signals against the market: (a) S&P 100 stock universe (2023-2024) and (b) S&P 500 stock universe (2024).
The attribution analysis (Table VI) reveals additional insights. With a 77–78% win rate across implementations, the system demonstrates remarkable consistency in signal precision. The positive alpha (17.6–18.9%) and elevated beta (1.24–1.27) of the S&P 500 portfolios suggest MarketSenseAI successfully identifies high-beta stocks with idiosyncratic upside potential. Furthermore, the stable monthly signal generation—35.1 ±7.95 buy signals for S&P 100 and 144.8 ±30.8 for S&P 500—indicates systematic selection rather than concentrated bets.
V-B Factor Analysis and Risk Decomposition
To elucidate the drivers of MarketSenseAI’s outperformance, we decompose portfolio returns (MS-Eq - S&P 100) using the Carhart four-factor [52] and Fama-French five-factor [53] models. Both models explain a substantial portion of return variance ($R^{2}=88.4\%$ and $85.4\%$, respectively), validating their applicability. Key findings are summarized in Table VII and discussed below.
V-B1 Market Exposure and Size Bias
MarketSenseAI exhibits near-neutral market exposure ($\beta=0.95$$0.96$). The negative SMB coefficients (0.13-0.13 to 0.22-0.22, $p<0.01$) reflect a tilt toward large-cap stocks, aligning with the S&P 100/500 universes [^2] SMB: size premium (small minus big capitalization stocks)..
V-B2 Value and Momentum Factors
Both models confirm consistent value exposure (HML = 0.08–0.11, p<0.01p<0.01) [^3] HML: value premium (high minus low book-to-market ratio stocks)., underscoring the Fundamentals Agent’s ability to identify undervalued equities through financial statement analysis. The Carhart model’s strong momentum loading (Mom = 0.18, p ¡ 0.01) highlights MarketSenseAI’s integration of price trends via the Dynamics Agent, a feature often absent in traditional fundamental models. This synergy between value and momentum aligns with the system’s architecture, where LLM-driven news sentiment and price dynamics reinforce fundamental insights.
V-B3 Profitability and Investment Factors
The five-factor model reveals insignificant loadings on profitability (RMW) and investment (CMA) [^4] RMW: profitability premium (robust minus weak firms); CMA: investment premium (conservative minus aggressive investment policies)., suggesting these factors play minimal roles in MarketSenseAI’s strategy. This suggests MarketSenseAI’s returns are not systematically driven by these traditional style factors. The system’s integration of multiple data sources may help identify alpha sources beyond conventional factor premiums.
V-B4 Alpha Generation and Unexplained Returns